What is displacement of an object in Simple Harmonic Motion when kinetic and potential energy are equal? please explain


1
Expert's answer
2021-07-05T08:43:27-0400

"E = E_k+E_p"

"E_k =\frac{mv^2}{2}"

"E_p =\frac{kx^2}{2}"

"\text{where x offset from equilibrium point position}"

"\text{let A oscillation amplitude,then for x = A:}"

"E=E'_k+E'_p"

"E'_k= 0 \text{ etc }v= 0"

"E_p =\frac{kA^2}{2}"

"E = \frac{kA^2}{2}"

"\text{if } E_p=E_k:"

"E_p+E_k= E;"

"2*E_p=\frac{kA^2}{2}"

"2*\frac{kx_1^2}{2}=\frac{kA^2}{2}"

"x_1= \sqrt{\frac{A^2}{2}}=0.71*A"

"\text{Answer: kinetic and potential energies are equal when shifted }"

"\text {by 0.71*A from the equilibrium point;}"

"\text{ A this is the Simple Harmonic Motion amplitude}"